Assume an goods and services market of an economy is characterized by the following

equations:

C = 0.8 (Y - T)

I = 800 -20r

Y = C + I + G

T = 1000

G = 1000

9. Consider for the moment the Keynesian Cross model. What will happen to the

GDP if G increases by 200? What is the multiplier?

10.Keep considering the Keynesian Cross model. What will happen to the GDP if T

increases by 200? What if both G and T increase by 200?

11.Derive a formula for the IS curve, showing Y as a function of r .

The money market for this economy is described by the equations:

(M/P) d = 0.4Y - 40r

M = 1200

P = 1

12.Derive a formula for the LM curve, showing Y as a function of r .

13.What are the short run values of Y and r ?

14.What are the short run values of Y and r if G increases by 200? What is the

multiplier? Is the value different from what you calculated for question 9? Explain

why it is different.

15.Derive a formula for the AD curve, showing Y as a function of P .

PLEASE ANWER QUESTION 12 TO 15
Answer #1

09 Multiplier =) 1-MPC 1-08 2 5 02 of G increases by 200 sY = multiplier X AG 5 x 200 lood ду will G, DP Increase by 1000 Q10of both increases DY ) -AG Simultaneo eously A-C 1-6 (-4) 200 + AT 1-C 5 X 200 + 1000 800 키 200 Overall GDP Increases by (200Qu Cane for IS Y = AD Y = Ct I to Y = 0.8 (Y- 2000) + 8co-dor + 10us y s 0.8 ya lo + goo 860 +86 – 200 thooo 4-0.8y = 1000 Xfor Equilibrium Ella 5000 1oor = 3000 floor oo 2oor eo I lo e =h 3000 + 100 (10) Y your CS Scanned with CamScannergly се 6, 9nceae) ву Is fan! у. Сf I f() У= o-s (у- соо) + Зоо-2» - 720 y= o - Уу - хво + yoo -)оэ 4 /20» у -68 y = ) 2оо - оAs Go Increases (powding out The auswer is different from @ because Is lume shifts Right word which raises enterest rate and50 y 30 t Y [ОО Р 100 sot 30 = 160 of y 100 P roo 24 50p + 30 100 as a (OE + jes) as Y assi f d asso у .P tasse 150 Р -Y CS S